it takes a plane traveling at a rate of 450 mph 25 minutes longer to travel a certain distance than it takes a second plane traveling at a rate of 500mph to travel the same distance. what is the distance.
all I realy need or want is help seeing the set up, after that im ok

Answer by mananth(16946)

let distance be y
time be x
distance = speed * time
y=450*(x-(25/60))
y=450*(x-5/12)
y= 450((12x-5)/12)
II plane
y= 500x
equate both equations
500x= 450(12x-5)/12
12*500x=450(12x-5)
6000x= 5400x-2250
600x=2250
x= 375 miles